A large rectangular movie screen in an IMAX theater has an area of 8,500 square feet. Find the dimensions of the screen if it is 15 feet longer than it is wide.

length       ft
width       ft
Answer #1

8lven xsy = 8500 - Y1y+I5) =8500 y?715Y-8500=0 7100y-s5Y-8500-0 %3D Length = 100 ft wicth = 85 ft ANS > =85 X= 100 %3D
Answer #2

SOLUTION :


L = W + 15 


So,


A = L * W = (W + 15) * W = W^2 + 15W 

=> 8500 = W^2 + 15W

=> W^2 + 15W - 8500 = 0

=> (W + 100)(W - 85) = 0

=> W = - 100, 85 


Negative value is meaningless.


So, W = 85 feet and L = 85 + 15 = 100 feet


So, 


 Length = 100 ft.


Width = 85 ft. 


(ANSWER).

answered by: Tulsiram Garg